We were told to factor this polynomial but I cannot figure out the Gcf.
4s + 6st - 14st

Answer by AnlytcPhil(1806)

We were told to factor this polynomial but I 
cannot figure out the Gcf. 

4s + 6st - 14st

Those last two terms are like terms, are you sure you
copied it right? Did you leave out an exponent? I'll
assume you didn't and I just won't combine like terms
till the last step:

Prime factor 4s, 6st and 14st

4s = 2·2·s
6st = 2·3·s·t
14st = 2·7·s·t

Line up the factors so that only like factors are 
lined up vertically, like this:  

 4s  = 2·2    ·s
 6st = 2  ·3  ·s·t
14st = 2    ·7·s·t
 
There are only two factors that appear in all three
rows, the ones I have colored red and blue:

 4s  = 2·2    ·s
 6st = 2  ·3  ·s·t
14st = 2    ·7·s·t 

So the GCF is 2·s 

4s + 6st - 14st =

2·2·s + 2·3·s·t - 2·7·s·t

So take out the red and blue factors, 
and put all the black ones inside
parentheses along with their signs:

2·s(2 + 3·t - 7·t)

or

2s(2 + 3t - 7t)

Notice that there are two like terms in the parentheses,
+3t and -7t which will give -4t

2s(2 - 4t)
 
Edwin